OpenStudy (anonymous):

What is the surface area of this prism? A. 184 cm2 B. 122 cm2 C. 240 cm2 D. 120 cm2

OpenStudy (anonymous):

@mathstudent55

OpenStudy (danjs):

It will be the area of 3 rectangles, plus area of 2 triangles

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

keep going

OpenStudy (anonymous):

stay and help me

OpenStudy (danjs):

base = 6 Height = 4 Area of Triangle = 1/2 * Base * Height

OpenStudy (anonymous):

ok

OpenStudy (danjs):

Then you have 2 (10 by 5) rectangles, and 1 (10 by 6) rectangle

OpenStudy (anonymous):

k

OpenStudy (danjs):

2 of the triangles areas will be Base x height = 6 * 4 = 24 cm^2

OpenStudy (anonymous):

ok

OpenStudy (danjs):

plus the rectangles area = length x width

OpenStudy (danjs):

24 cm^2 + 100 cm^2 + 60 cm^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (danjs):

the total surface area is (24 + 100 + 60) cm^2

OpenStudy (danjs):

welcome

OpenStudy (anonymous):

here

OpenStudy (anonymous):

so the answer is 184

OpenStudy (danjs):

yea

OpenStudy (danjs):

cm^2

OpenStudy (anonymous):

can you help me a few more